EV003

PI: Venturi
Title: 3C317

Session: February 1996 (session 1)
Frequency: 18 cm
Mode: E
Obs time (UT): 44 EU 00:00(14/02)-11:00(14/02)
Stations: SLNJWShUr+MERLIN
Processor: Bonn
EU Mk3 Tapes/St.: 2

Comment


To give feedback on this experiment, click here

Experiment Feedback

Onsala (John Conway at Mon Feb 19 08:19:25 EST 1996) Minor failures
First experiment after repair of Capstan motor and Capstan. Phase-cal unit was initially on but later switched off because it was increasing Tsys in some channels.

Urumqi (Lee Suqin at Wed Feb 21 09:12:04 EST 1996) Minor failures
Cal temprature is 103k and the system temprature is about 210k. It is cloudy. There is two scan have not record in full. There is some error in schadule scan numbers. Observation is OK.

Jodrell (Dave Henstock at Thu Feb 22 06:44:28 EST 1996) Minor failures
RFI in observing band throughout experiment. Particularly bad in VCs 9,10. Due to RFI, Tsys measurements only crudely calibrated. Note schedule had wrong telescope (should be Lovell=JODRELL1, not JODRELL2) so beware when correlating. Specific Problems: Several times telescope late on source due to slew. Day 45 00:00-00:25 and 03:11-03:41 and 06:30-06:43 and 08:12-08:32 Field System crashed. Scans/passes between these timnes will have been lost. Tsys is in Jy, need to multiply measured Tsys by 1.23 for better calibration.

Cambridge (Dave Henstock at Thu Feb 22 06:45:31 EST 1996) Success
No problems. Note that there are no Tsys measurements from Cambridge. Tsys is 212 Jy (from EVN status).

Medicina (Venturi Tiziana at Thu Feb 22 09:58:26 EST 1996) Success
Experiment carried out successfully.

Noto (Corrado Trigilio at Sat Feb 24 06:26:31 EST 1996) Success
No problems reported. Corrected Tsys values will be provvided.

Westerbork (foley at Mon Feb 26 11:45:43 EST 1996) Success
Note formatter reset before observation

Shanghai (Huang Xinyong at Fri Mar 1 04:32:31 EST 1996) Minor failures
No problems with observations. Weather fine. Clock offset = -0.85 usec. Observations were in LCP with phase cal signal present.

Urumqi (Leesuqin at Thu Mar 7 00:20:12 EST 1996) Minor failures

Calibration (Fredrik T Rantakyro at Thu Sep 26 10:10:39 EDT 1996) Minor failures
An ANTAB formatted file for the experiment can be found on VLBEER